A collection of nickels, dimes, and quarters consist of
60 coins with a total of
$6.00. If there are
2 times as many dimes as quarters, find the number of each type of coins.

Answer by Alan3354(69443)

A collection of nickels, dimes, and quarters consist of
60 coins with a total of
$6.00. If there are
2 times as many dimes as quarters, find the number of each type of coins.
----------
n + d + q = 60
d = 2q
5n + 10d + 25q = 600
------
n + 3q = 60
5n + 45q = 600 --> n + 9q = 120
---
n + 9q = 120
n + 3q = 60
------------------ Subtract
6q = 60
q = 10 quarters
d = 20 dimes
n = 30 nickels